# Physics-Informed Teleconnection-Aware Transformer for Global Subseasonal-to-Seasonal Forecasting

### v2

@@ 4 Experiments
We implement all models using PyTorch Lightning and train our direct training baselines with consistent hyperparameters: batch size 16, hidden dimension 256, learning rate 0.01, $\color[rgb]{0.171875,0.2421875,0.3125}\gamma=0.1$, and $\color[rgb]{0.171875,0.2421875,0.3125}\lambda=0.2$. For simplicity, we initialize $\color[rgb]{0.171875,0.2421875,0.3125}\mathbf{P}_{j}$ as a trainable vector that enables to learn the teleconnection patterns from data directly. All training was conducted on 4 GeForce RTX A40 GPUs. ClimaX and CirT were trained using same configuration as TelePiT, while pre-trained versions of FourCastNetV2, Pangu-Weather, and GraphCast were accessed through ECMWF’s API [^1] https://github.com/ecmwf-lab/ai-models. Inference for these pre-trained models was performed on NVIDIA A800 80G GPUs. It should be noted that although FourCastNetV2 and Pangu-Weather reportedly support $\color[rgb]{0.171875,0.2421875,0.3125}t2m$ predictions, this capability was not available in the ECMWF-provided implementations. Additionally, GraphCast encountered out-of-memory when attempting inference for the extended Weeks 5-6 forecast period.
Figure 2: RMSE comparison on variable $\color[rgb]{0.171875,0.2421875,0.3125}z$, $\color[rgb]{0.171875,0.2421875,0.3125}t$, $\color[rgb]{0.171875,0.2421875,0.3125}u$, and $\color[rgb]{0.171875,0.2421875,0.3125}v$ of different pressure levels.
@@
Due to space constraints, we present results for key variables and primary metrics in the main text. More additional experiments can be found in the Appendix B.
4.1 Overall Performance
Compared with data-driven models
Table 1 and Figure 2 present a comprehensive comparison of TelePiT against state-of-the-art data-driven S2S forecasting models. Our model demonstrates consistent and substantial improvements across all meteorological variables and forecast horizons. For geopotential heights ($\color[rgb]{0.171875,0.2421875,0.3125}z500$, $\color[rgb]{0.171875,0.2421875,0.3125}z850$), TelePiT reduces RMSE by 9.5% and 8.6% respectively compared to the previous best model (CirT) at Weeks 3-4, while simultaneously achieving higher ACC values. This improvement is particularly noteworthy given that these upper-atmospheric variables govern large-scale circulation patterns crucial for extended-range forecasting.
The most striking performance advantage appears in near-surface variables. For 2-meter temperature ($\color[rgb]{0.171875,0.2421875,0.3125}t2m$), TelePiT achieves an RMSE of 12.057K at Weeks 3-4, representing a remarkable 57.7% reduction compared to CirT (28.526K) and 80.2% reduction compared to ClimaX (60.894K). This exceptional improvement for surface temperature, the variable with the greatest socioeconomic relevance, demonstrates TelePiT’s particular strength in capturing complex land-atmosphere interactions.
Compared with numerical models
Most notably, TelePiT demonstrates exceptional performance for temperature ($\color[rgb]{0.171875,0.2421875,0.3125}t$) forecasts throughout the troposphere and lower stratosphere, with substantial improvements over physics-based models that have benefited from decades of scientific development and operational refinement. This advantage is particularly pronounced at upper levels (10-200 hPa), where traditional models struggle to maintain forecast skill beyond week 2. Critically, while physics-based models show substantial skill degradation between Weeks 3-4 and Weeks 5-6 (evidenced by darker colors in the right panels), TelePiT maintains more consistent performance across both forecast horizons.
4.2 Ablation Study
The ablation study presented in Table 3 systematically evaluates the contribution of each component in our proposed TelePiT architecture. Spherical Harmonic Embedding (SHE) provides the most critical contribution, particularly for $\color[rgb]{0.171875,0.2421875,0.3125}t2m$ prediction where its removal increases RMSE from 12.057 to 27.206 at Weeks 3-4. This substantial degradation demonstrates that effectively encoding Earth’s spherical geometry is fundamental for capturing planetary-scale circulation patterns. Wavelet Decomposition (WD) plays an essential role in separating signals across temporal frequencies. Its removal particularly impacts geopotential height predictions ($\color[rgb]{0.171875,0.2421875,0.3125}z500$, $\color[rgb]{0.171875,0.2421875,0.3125}z850$), with RMSE increasing by 7-8%. This confirms our hypothesis that atmospheric dynamics operate simultaneously across multiple time scales, from fast-moving synoptic systems to slow-evolving planetary waves, requiring explicit multi-scale processing for accurate prediction.
Physics-Informed ODE shows more subtle but consistent contributions, particularly for near-surface wind components. Unlike purely data-driven approaches, the integration of simplified atmospheric dynamics through neural ODEs provides physical consistency that enhances model generalization, especially for variables with complex boundary-layer interactions. Teleconnection Attention (TA) becomes increasingly valuable at longer lead times, with its removal causing notable degradation at Weeks 5-6. This aligns with meteorological understanding that as initial condition influence diminishes, large-scale teleconnections increasingly govern predictability.
Table 3: Ablation study of different TelePiT model variants.
@@ 4.2 Ablation Study
Figure 3: The global RMSE distribution of $\color[rgb]{0.171875,0.2421875,0.3125}t850$ with lead times weeks 3-4 in the testing set.
4.3 Empirical Analysis
Global visualization
@@ 4.3 Empirical Analysis
Figure 3 visualizes the global distribution of RMSE for $\color[rgb]{0.171875,0.2421875,0.3125}t850$ at Weeks 3-4. TelePiT demonstrates remarkably consistent low error (deep blue) across nearly all geographical regions, including traditionally challenging areas like continental interiors and the polar regions. This spatial analysis confirms that TelePiT’s architectural innovations, especially its spherical harmonic embedding and teleconnection attention mechanisms, enable it to effectively capture both tropical atmospheric dynamics and mid-to-high latitude variability, areas where existing models often struggle due to fundamentally different governing physical processes.
Forecasting visualization
@@
Figure 4 presents the $\color[rgb]{0.171875,0.2421875,0.3125}t850$ forecast performance on each day. TelePiT demonstrates consistently lower RMSE values throughout the annual cycle for both Weeks 3-4 and Weeks 5-6 forecasts, maintaining this advantage across all seasonal transitions. The performance gap widens even further in the more challenging Weeks 5-6 timeframe, where TelePiT maintains relatively stable error metrics while baselines show increased variability and deterioration. This consistent temporal superiority complements the spatial advantages described in the global visualization analysis, suggesting that TelePiT’s architectural design effectively captures both the spatial and temporal aspects of atmospheric dynamics critical for extended-range forecasting.
Figure 4: Forecasting Performance of $\color[rgb]{0.171875,0.2421875,0.3125}t850$ on each day in the testing set.
5 Related Works
@@
S2S forecasting has traditionally relied on sophisticated numerical weather prediction (NWP) systems vitart2017subseasonal ; white2017potential , which integrate atmospheric and oceanic physics through coupled dynamical models. These systems, exemplified by ECMWF molteni1996ecmwf , UKMO williams2015met , NCEP saha2014ncep , and CMA wu2019beijing , have gradually extended forecast horizons through enhanced initialization techniques and ensemble methods vitart2018sub . However, these methods remain computationally intensive and often struggle to fully capture complex teleconnection patterns that significantly influence predictability beyond the 2-week deterministic limit bauer2015quiet ; robertson2020subseasonal . Recent advancements utilizing empirical methods to extract large-scale modes of variability cohen2019s2s ; merryfield2020current have shown promise but lack physical dynamics across scales.
@@
Deep learning has revolutionized climate prediction. Foundational work by Rasp et al. rasp2020weatherbench established benchmarks for data-driven atmospheric forecasting, while subsequent architectures including FourCastNetV2 kurth2023fourcastnet , Pangu-Weather bi2023accurate , GraphCast lam2023learning , and ClimaX nguyen2023climax demonstrated competitive performance against operational NWP systems for medium-range forecasting. Regional S2S forecasting has seen notable advances through specialized models like SubseasonalClimateUSA mouatadid2023subseasonalclimateusa , which focuses on North American climate predictions, and approaches by He et al. he2022learning and Hwang et al. hwang2019improving that target localized weather patterns. However, these regional models often overlook crucial global teleconnections. Recent global S2S efforts include FuXi-S2S chen2024machine , which employs an encoder-decoder framework for global precipitation forecasting, and CirT liucirt , which incorporates spherical geometry. Nevertheless, these approaches lack explicit incorporation of physical constraints and multi-scale atmospheric processes, limiting their ability to maintain skill at extended lead times nathaniel2024chaosbench .
@@
Physics-informed neural networks (PINNs) have shown increasing promise for integrating physical knowledge into data-driven models raissi2019physics ; karniadakis2021physics . In climate science, Reichstein et al. reichstein2019deep and Willard et al. willard2022integrating established frameworks for integrating scientific knowledge with machine learning for Earth system modeling. Recent work on climate modeling with neural advection-diffusion equations choi2023climate has demonstrated how neural networks can parameterize complex climate dynamics while preserving physical principles. Complementary approaches in machine learning-accelerated computational fluid dynamics kochkov2021machine have shown that neural networks can significantly enhance traditional solvers while maintaining physical consistency. Unlike these works, which typically focus on individual atmospheric processes or fixed spatial scales, TelePiT uniquely combines learnable multi-scale decomposition with physics-informed ODEs that adaptively capture scale-dependent dynamics. For teleconnection modeling, previous efforts cohen2021linking have explored statistical approaches, while Ham et al. ham2019deep pioneered deep learning for El Niño prediction. TelePiT advances beyond these approaches by seamlessly integrating scale-aware physics with teleconnection-aware attention mechanisms, explicitly capturing the cross-scale interactions and global patterns that are crucial for S2S predictability while preserving Earth’s spherical geometry in a unified end-to-end architecture.
6 Conclusion
@@
We presented TelePiT, a novel deep learning architecture for global S2S forecasting that integrates physical principles with data-driven learning. By combining Spherical Harmonic Embedding, Multi-Scale Physics-Informed ODE, and Teleconnection-Aware Transformer, we explicitly model the fundamental processes governing S2S predictability. Our evaluation shows that TelePiT consistently outperforms both state-of-the-art data-driven models and operational numerical systems across all meteorological variables and forecast horizons. The most significant improvements occur in near-surface variables, particularly 2-meter temperature, with over 57% RMSE reduction compared to previous models. Ablation studies confirm each component’s meaningful contribution, with their combined effect becoming more pronounced at extended forecasting ranges. This advances climate resilience applications in agriculture, energy, and disaster preparedness. Future work may incorporate additional Earth system components, further bridging the gap between weather and climate prediction.
@@
  • A. Theoretical Analysis A
    • A.1. Spectral Representation on the Sphere A.1
    • A.2. Multi-Scale Analysis and Function Approximation A.2
    • A.3. Physics-Informed Neural ODEs A.3
    • A.4. Teleconnection Representation Theory A.4
    • A.5. Complexity and Generalization Analysis A.5
  • B. Additional Experimental Details B
    • B.1. Evaluation Metric B.1
    • B.2. Baselines B.2
    • B.3. Additional Overall Performance B.3
    • B.4. Additional Ablation Study B.4
    • B.5. Additional Empirical Analysis B.5
    • B.6. Robustness Study B.6
    • B.7. Parameter Sensitivity Analysis B.7
Overview
The main paper presents core results including overall performance of RMSE and ACC metrics for key variables ($\color[rgb]{0.171875,0.2421875,0.3125}z500,z850,t500,t850,t2m,u10,v10$) [Table 1], SpecDiv for physics-based evaluation [Table 2], heatmap comparisons across pressure levels [Figure 2], ablation study on RMSE [Table 3], global visualization of RMSE distribution for $\color[rgb]{0.171875,0.2421875,0.3125}t850$ [Figure 3], and daily forecasting performance visualization for $\color[rgb]{0.171875,0.2421875,0.3125}t850$ [Figure 4]. Figure 5 offers an overview of how experiments are distributed between the main paper and appendix.
This appendix significantly extends these analyses with:
  1. 1.
  2. 2.
  3. 3.
  4. 4.
    Comprehensive Variable Coverage: We present detailed results for all pressure levels and variables, including comprehensive heatmaps for ACC [Figure 6], MS-SSIM [Figure 7], SpecDiv [Figure 8], and SpecRes [Figure 9] metrics.
  5. 5.
  6. 6.
    Additional Global Visualizations: We expand global visualizations of RMSE distribution to cover key variables ($\color[rgb]{0.171875,0.2421875,0.3125}t850$, $\color[rgb]{0.171875,0.2421875,0.3125}t500$, $\color[rgb]{0.171875,0.2421875,0.3125}z850$, $\color[rgb]{0.171875,0.2421875,0.3125}z500$) for both forecast horizons (Weeks 3-4 and Weeks 5-6) [Figure 10,11,12,13,14,15,16].
  7. 7.
    Daily Forecasting Performance: We include additional daily forecasting performance (RMSE) for $\color[rgb]{0.171875,0.2421875,0.3125}t500$ [Figure 17], $\color[rgb]{0.171875,0.2421875,0.3125}z850$ [Figure 18] and $\color[rgb]{0.171875,0.2421875,0.3125}z500$ [Figure 19] in 2018, demonstrating TelePiT’s consistent advantages across seasonal transitions.
  8. 8.
  9. 9.
    Parameter Sensitivity Analysis: We analyze the impact of the teleconnection coefficient $\color[rgb]{0.171875,0.2421875,0.3125}\lambda$ on model performance, revealing optimal parameter settings and validating architectural design choices [Figure 20].
(a) Distribution of overall performance for key variables $\color[rgb]{0.171875,0.2421875,0.3125}z500$, $\color[rgb]{0.171875,0.2421875,0.3125}z850$, $\color[rgb]{0.171875,0.2421875,0.3125}t500$, $\color[rgb]{0.171875,0.2421875,0.3125}t850$, $\color[rgb]{0.171875,0.2421875,0.3125}t2m$, $\color[rgb]{0.171875,0.2421875,0.3125}u10$, $\color[rgb]{0.171875,0.2421875,0.3125}v10$.
(b) Distribution of heatmap visualizations for variables ($\color[rgb]{0.171875,0.2421875,0.3125}z$, $\color[rgb]{0.171875,0.2421875,0.3125}t$, $\color[rgb]{0.171875,0.2421875,0.3125}u$, $\color[rgb]{0.171875,0.2421875,0.3125}v$) across all pressure levels.
@@
Figure 5: Experiments distribution between main paper and appendix. Note that GraphCast (GC) results are unavailable at Weeks 5-6 due to out-of-memory issues during inference. UKMO’s global RMSE visualizations are not shown due to significant missing values in their inference results.

### v3

@@ 4. Experiments
Experimental setup. We implement all models using PyTorch Lightning and train our direct training baselines with consistent hyperparameters: batch size 16, hidden dimension 256, learning rate 0.01, $\gamma=0.1$, and $\lambda=0.2$. For simplicity, we initialize $\mathbf{P}_{j}$ as a trainable vector that enables to learn the teleconnection patterns from data directly. All training was conducted on 4 GeForce RTX A40 GPUs. ClimaX and CirT were trained using same configuration as TelePiT, while pre-trained versions of FourCastNetV2, Pangu-Weather, and GraphCast were accessed through ECMWF’s API [^1] https://github.com/ecmwf-lab/ai-models. Inference for these pre-trained models was performed on NVIDIA A800 80G GPUs. It should be noted that although FourCastNetV2 and Pangu-Weather reportedly support $t2m$ predictions, this capability was not available in the ECMWF-provided implementations. Additionally, GraphCast encountered out-of-memory when attempting inference for Weeks 5-6 forecast period.
@@
Due to space constraints, we present results for key variables and primary metrics in the main text. More additional experiments and model complexity analysis can be found in the Appendix B.
Table 4. Ablation study of different TelePiT model variants.
@@
4.1. Overall Performance
4.1.1. Compared with data-driven models.
Table 1 and Figure 2 present a comprehensive comparison of TelePiT against state-of-the-art data-driven S2S forecasting models. Our model demonstrates consistent and substantial improvements across all meteorological variables and forecast horizons. For geopotential heights ($z500$, $z850$), TelePiT reduces RMSE by 9.5% and 8.6% respectively compared to the CirT at Weeks 3-4, while simultaneously achieving higher ACC values. This improvement is particularly noteworthy given that these upper-atmospheric variables govern large-scale circulation patterns crucial for extended-range forecasting.
@@ 4.1. Overall Performance
Figure 3. The global RMSE distribution of $t850$ with lead times weeks 3-4 in the testing set.
4.1.2. Compared with numerical models.
Most notably, TelePiT demonstrates exceptional performance for temperature ($t$) forecasts throughout the troposphere and lower stratosphere, with substantial improvements over physics-based models that have benefited from decades of scientific development and operational refinement. This advantage is particularly pronounced at upper levels (10-200 hPa), where traditional models struggle to maintain forecast skill beyond week 2. Critically, while physics-based models show substantial skill degradation between Weeks 3-4 and Weeks 5-6 (evidenced by darker colors in the right panels), TelePiT maintains more consistent performance across both forecast horizons.
4.2. Ablation Study
The ablation study presented in Table 4 systematically evaluates the contribution of each component in our proposed TelePiT architecture. Spherical Harmonic Embedding (SHE) provides the most critical contribution, particularly for $t2m$ prediction where its removal increases RMSE from 12.057 to 27.206 at Weeks 3-4. This substantial degradation demonstrates that effectively encoding Earth’s spherical geometry is fundamental for capturing planetary-scale circulation patterns. Wavelet Decomposition (WD) plays an essential role in separating signals across temporal frequencies. Its removal impacts geopotential height predictions ($z500$, $z850$), with RMSE increasing by 7-8%. This confirms our hypothesis that atmospheric dynamics operate simultaneously across multiple time scales, from fast-moving synoptic systems to slow-evolving planetary waves, requiring explicit multi-scale processing for accurate prediction.
Physics-Informed ODE shows more subtle but consistent contributions. Unlike purely data-driven approaches, the integration of simplified atmospheric dynamics through neural ODEs provides physical consistency that enhances model generalization, especially for variables with complex boundary-layer interactions. Teleconnection Attention (TA) becomes increasingly valuable at longer lead times, with its removal causing notable degradation at Weeks 5-6. This aligns with meteorological understanding that as initial condition influence diminishes, large-scale teleconnections increasingly govern predictability.
Figure 4. Forecasting Performance of $t850$ on each day in the testing set.
4.3. Empirical Analysis
@@
Global visualization. Figure 3 visualizes the global distribution of RMSE for $t850$ at Weeks 3-4. TelePiT demonstrates remarkably consistent low error (deep blue) across nearly all geographical regions, including traditionally challenging areas like continental interiors and the polar regions. This spatial analysis confirms that TelePiT’s architectural innovations, especially its spherical harmonic embedding and teleconnection attention mechanisms, enable it to effectively capture both tropical atmospheric dynamics and mid-to-high latitude variability, areas where existing models often struggle due to fundamentally different governing physical processes.
@@
Forecasting visualization. Figure 4 presents the $t850$ forecast performance on each day. TelePiT demonstrates consistently lower RMSE values throughout the annual cycle for both Weeks 3-4 and Weeks 5-6 forecasts, maintaining this advantage across all seasonal transitions. The performance gap widens even further in the more challenging Weeks 5-6 timeframe, where TelePiT maintains relatively stable error metrics while baselines show increased variability and deterioration. This consistent temporal superiority complements the spatial advantages described in the global visualization analysis, suggesting that TelePiT’s architectural design effectively captures both the spatial and temporal aspects of atmospheric dynamics critical for extended-range forecasting.
5. Related Works
@@
S2S forecasting has traditionally relied on sophisticated numerical weather prediction (NWP) systems (Vitart et al., 2017; White et al., 2017), which integrate atmospheric and oceanic physics through coupled dynamical models. These systems, exemplified by ECMWF (Molteni et al., 1996), UKMO (Williams et al., 2015), NCEP (Saha et al., 2014), and CMA (Wu et al., 2019), have gradually extended forecast horizons through enhanced initialization techniques and ensemble methods (Vitart and Robertson, 2018). However, these methods remain computationally intensive and often struggle to fully capture complex teleconnection patterns that significantly influence predictability beyond the 2-week deterministic limit (Bauer et al., 2015; Robertson et al., 2020). Recent advancements utilizing empirical methods to extract large-scale modes of variability (Cohen et al., 2019; Merryfield et al., 2020) have shown promise but lack physical dynamics across scales.
@@
Deep learning has revolutionized climate prediction. Foundational work by Rasp et al. (Rasp et al., 2020) established benchmarks for data-driven atmospheric forecasting, while subsequent architectures including FourCastNetV2 (Kurth et al., 2023), Pangu-Weather (Bi et al., 2023), GraphCast (Lam et al., 2023), and ClimaX (Nguyen et al., 2023) demonstrated competitive performance against operational NWP systems for medium-range forecasting. Regional S2S forecasting has seen notable advances through specialized models like SubseasonalClimateUSA (Mouatadid et al., 2023), which focuses on North American climate predictions, and approaches by He et al. (He et al., 2022) and Hwang et al. (Hwang et al., 2019) that target localized weather patterns. However, these regional models often overlook crucial global teleconnections. Recent global S2S efforts include FuXi-S2S (Chen et al., 2024), which employs an encoder-decoder framework for global precipitation forecasting, and CirT (Liu et al., 2025), which incorporates spherical geometry. Nevertheless, these approaches lack explicit incorporation of physical constraints and multi-scale atmospheric processes, limiting their ability to maintain skill at extended lead times (Nathaniel et al., 2024).
@@
Physics-informed neural networks (PINNs) have shown increasing promise for integrating physical knowledge into data-driven models (Raissi et al., 2019; Karniadakis et al., 2021). In climate science, Reichstein et al. (Reichstein et al., 2019) and Willard et al. (Willard et al., 2022) established frameworks for integrating scientific knowledge with machine learning for Earth system modeling. Recent work on climate modeling with neural advection-diffusion equations (Choi et al., 2023) has demonstrated how neural networks can parameterize complex climate dynamics while preserving physical principles. Complementary approaches in machine learning-accelerated computational fluid dynamics (Kochkov et al., 2021) have shown that neural networks can significantly enhance traditional solvers while maintaining physical consistency. Unlike these works, which typically focus on individual atmospheric processes or fixed spatial scales, TelePiT uniquely combines learnable multi-scale decomposition with physics-informed ODEs that adaptively capture scale-dependent dynamics. For teleconnection modeling, previous efforts (Cohen et al., 2021) have explored statistical approaches, while Ham et al. (Ham et al., 2019) pioneered deep learning for El Niño prediction. TelePiT advances beyond these approaches by seamlessly integrating scale-aware physics with teleconnection-aware attention mechanisms, explicitly capturing the cross-scale interactions and global patterns that are crucial for S2S predictability while preserving Earth’s spherical geometry in a unified end-to-end architecture.
6. Conclusion
@@
We presented TelePiT, a novel deep learning architecture for global S2S forecasting that integrates physical principles with data-driven learning. By combining Spherical Harmonic Embedding, Multi-Scale Physics-Informed ODE, and Teleconnection-Aware Transformer, we explicitly model the fundamental processes governing S2S predictability. Our evaluation shows that TelePiT consistently outperforms both state-of-the-art data-driven models and operational numerical systems across all meteorological variables and forecast horizons. The most significant improvements occur in near-surface variables, particularly 2-meter temperature, with over 57% RMSE reduction compared to previous models. Ablation studies confirm each component’s meaningful contribution, with their combined effect becoming more pronounced at extended forecasting ranges. This advances climate resilience applications in agriculture, energy, and disaster preparedness. Future research directions include incorporating additional Earth system components and specialized datasets such as the Arctic System Reanalysis (ASR) to enhance regional performance, ultimately bridging the gap between weather and climate prediction scales.
@@
  • A. Theoretical Analysis .A
    • A.1. Spectral Representation on the Sphere .A.1
    • A.2. Multi-Scale Analysis and Function Approximation .A.2
    • A.3. Physics-Informed Neural ODEs .A.3
    • A.4. Teleconnection Representation Theory .A.4
    • A.5. Complexity and Generalization Analysis .A.5
  • B. Additional Experimental Details .B
    • B.1. Evaluation Metric .B.1
    • B.2. Baselines .B.2
    • B.3. Model Complexity .B.3
    • B.4. Additional Overall Performance .B.4
    • B.5. Additional Ablation Study .B.5
    • B.6. Additional Empirical Analysis .B.6
    • B.7. Robustness Study .B.7
    • B.8. Parameter Sensitivity Analysis .B.8
(a) Distribution of overall performance for key variables $z500$, $z850$, $t500$, $t850$, $t2m$, $u10$, $v10$.
(b) Distribution of heatmap visualizations for variables ($z$, $t$, $u$, $v$) across all pressure levels.
@@
Figure 5. Experiments distribution between main paper and appendix. Note that GraphCast (GC) results are unavailable at Weeks 5-6 due to out-of-memory issues during inference. UKMO’s global RMSE visualizations are not shown due to significant missing values in their inference results.
Overview. The main paper presents core results including overall performance of RMSE and ACC metrics for key variables ($z500$, $z850,t500,t850,t2m,u10,v10$) [Table 1], SpecDiv for physics-based evaluation [Table 2], heatmap comparisons across pressure levels [Figure 2], ablation study on RMSE [Table 4], global visualization of RMSE distribution for $t850$ [Figure 3], and daily forecasting performance visualization for $t850$ [Figure 4]. Figure 5 offers an overview of how experiments are distributed between the main paper and appendix.
This appendix significantly extends these analyses with:
  1. (1)
  2. (2)
    Model Complexity Analysis: We provide a comprehensive comparison of computational requirements across several key baselines [Table 5].
  3. (3)
  4. (4)
  5. (5)
    Comprehensive Variable Coverage: We present detailed results for all pressure levels and variables, including comprehensive heatmaps for ACC [Figure 6], MS-SSIM [Figure 7], SpecDiv [Figure 8], SpecRes [Figure 9] metrics.
  6. (6)
  7. (7)
    Additional Global Visualizations: We expand global visualizations of RMSE distribution to cover key variables ($t850$, $t500$, $z850$, $z500$) for both forecast horizons (Weeks 3-4 and Weeks 5-6) [Figure 10,11,12,13,14,15,16].
  8. (8)
    Daily Forecasting Performance: We include additional daily forecasting performance (RMSE) for $t500$ [Figure 17], $z850$ [Figure 18] and $z500$ [Figure 19] in 2018, demonstrating TelePiT’s consistent advantages across seasonal transitions.
  9. (9)
  10. (10)
    Parameter Sensitivity Analysis: We analyze the impact of the teleconnection coefficient $\lambda$ on model performance, revealing optimal parameter settings and validating architectural design choices [Figure 20].